# MapReduce LoRA: Advancing the Pareto Front in Multi-Preference Optimization for Generative Models

### v2

@@ C.2 Base Models and Reward Models
Reward Models Links
GenEval [GenEval_NeurIPS'23] https://github.com/djghosh13/geneval
PickScore [PickScore_NeurIPS'23] https://huggingface.co/yuvalkirstain/PickScore_v1
OCR [paddleocr_arxiv'25] https://github.com/PaddlePaddle/PaddleOCR
VQAScore [VQAScore_ECCV24] https://github.com/linzhiqiu/t2v_metrics
MPS [MPS_CVPR'24] https://github.com/Kwai-Kolors/MPS
VILA [VILA_CVPR'23] https://github.com/google-research/google-research/tree/master/vila
VQ and MQ [videoalign_arxiv25] https://github.com/KlingTeam/VideoAlign/tree/main
@@ Appendix D More Quantitative Results
Table 5: 3D merging results on Text-to-Image tasks: SD 3.5 M [sd35m_hf] and FLUX.1-dev [flux24].
@@
Table 6: 2D merging results on Text-to-Image tasks, SD 3.5 M [sd35m_hf] and FLUX.1-dev [flux24], and Text-to-Video task, HunyuanVideo [hunyuanvideo_arxiv'25].
@@
To compare naive multi-objective RL with data mixing (MORL-D) against MapReduce LoRA, we train MORL-D under ten {GenEval:PickScore:OCR} sampling ratios—{1:1:1}, {1:2:3}, {1:3:2}, {2:1:3}, {2:3:1}, {3:1:2}, {3:2:1}, {1:1:4}, {1:4:1}, {4:1:1}—using the same number of training steps as MapReduce LoRA on SD 3.5 M [sd35m_hf]. Fig. 10 compares Pareto fronts for Rewarded Soup [reward_soup_nips'23], MORL-D, and MapReduce LoRA (left: 3D; right: 2D projections). In 3D, MORL-D is confined to a small region across the three rewards. For readability, we also show 2D projections of the 3D front; because MORL-D is an a priori method, we cannot fix one reward to 0 as in Fig. 9, so the right panels are projections rather than true 2D Pareto fronts. Although MORL-D is competitive on GenEval and OCR, conflicting objectives reduce visual quality (PickScore; see Fig. 15), yielding only limited PickScore gains.
Figure 10: Merging performance comparison across Rewarded Soup [reward_soup_nips'23], MORL-D, and MapReduce LoRA. Left: 3D Pareto-front comparison. Right: 2D projections of the 3D Pareto front (for readability), which are not 2D Pareto fronts. Unlike Fig. 9, these are projections rather than a 2D merge with the third reward fixed to 0. MORL-D performance is confined to a small region across the three rewards and yields only limited improvement on PickScore.
@@ E.1 Text-to-Image qualitative results
Figs 11, 12 and 13 demonstrate the visual comparison of two rewards with different merging ratios. Figs 14 and 15 demonstrate the visual comparison of all methods, including Flow-GRPO [flow_grpo_arxiv'25] tuned on GenEval [GenEval_NeurIPS'23], PickScore [PickScore_NeurIPS'23], OCR [paddleocr_arxiv'25], Rewarded Soup [reward_soup_nips'23], MORL-D, MORL-DR and MapReduce LoRA. PickScore and OCR guidance cause strong reward overfitting: the former drives the model toward a narrow, high-scoring aesthetic style, while the latter encourages increasingly large and prominent text regardless of overall visual harmony or contextual appropriateness. Our proposed MapReduce LoRA mitigates these issues by iteratively merging per-reward experts, enabling the model to discover a more balanced (and often near-optimal) preference that harmonizes aesthetic fidelity with text readability.
@@ E.2 Text-to-Video qualitative results
Fig. 16 demonstrates the visual comparison of HunyuanVideo [hunyuanvideo_arxiv'25], Rewarded Soup [reward_soup_nips'23] and our proposed MapReduce LoRA. Also, Figs. 17 and 18 demonstrate the visual performance of merging progress that advances the visual quality, motion quality, and prompt alignment.
Figure 16: Visual comparison of HunyuanVideo [hunyuanvideo_arxiv'25], Rewarded Soup [reward_soup_nips'23], and MapReduce LoRA (Ours). In the upper case, MapReduce LoRA more faithfully renders the intended motion and writing action described in the prompt. In the lower case, MapReduce LoRA better adheres to the specified drone-movement control, producing a trajectory that aligns with the prompt.

### v3

@@ Appendix D More Quantitative Results
Table 5: 3D merging results on Text-to-Image tasks: SD 3.5 M [3] and FLUX.1-dev [17].
@@
Table 6: 2D merging results on Text-to-Image tasks, SD 3.5 M [3] and FLUX.1-dev [17], and Text-to-Video task, HunyuanVideo [16].
@@
To compare naive multi-objective RL with data mixing (MORL-D) against MapReduce LoRA, we train MORL-D under ten {GenEval:PickScore:OCR} sampling ratios—{1:1:1}, {1:2:3}, {1:3:2}, {2:1:3}, {2:3:1}, {3:1:2}, {3:2:1}, {1:1:4}, {1:4:1}, {4:1:1}—using the same number of training steps as MapReduce LoRA on SD 3.5 M [3]. Fig. 10 compares Pareto fronts for Rewarded Soup [33], MORL-D, and MapReduce LoRA (left: 3D; right: 2D projections). In 3D, MORL-D is confined to a small region across the three rewards. For readability, we also show 2D projections of the 3D front; because MORL-D is an a priori method, we cannot fix one reward to 0 as in Fig. 9, so the right panels are projections rather than true 2D Pareto fronts. Although MORL-D is competitive on GenEval and OCR, conflicting objectives reduce visual quality (PickScore; see Fig. 15), yielding only limited PickScore gains.
Figure 10: Merging performance comparison across Rewarded Soup [33], MORL-D, and MapReduce LoRA. Left: 3D Pareto-front comparison. Right: 2D projections of the 3D Pareto front (for readability), which are not 2D Pareto fronts. Unlike Fig. 9, these are projections rather than a 2D merge with the third reward fixed to 0. MORL-D performance is confined to a small region across the three rewards and yields only limited improvement on PickScore.
@@ E.1 Text-to-Image qualitative results
Figs 11, 12 and 13 demonstrate the visual comparison of two rewards with different merging ratios. Figs 14 and 15 demonstrate the visual comparison of all methods, including Flow-GRPO [23] tuned on GenEval [11], PickScore [15], OCR [6], Rewarded Soup [33], MORL-D, MORL-DR and MapReduce LoRA. PickScore and OCR guidance cause strong reward overfitting: the former drives the model toward a narrow, high-scoring aesthetic style, while the latter encourages increasingly large and prominent text regardless of overall visual harmony or contextual appropriateness. Our proposed MapReduce LoRA mitigates these issues by iteratively merging per-reward experts, enabling the model to discover a more balanced (and often near-optimal) preference that harmonizes aesthetic fidelity with text readability.
@@ E.2 Text-to-Video qualitative results
Fig. 16 demonstrates the visual comparison of HunyuanVideo [16], Rewarded Soup [33] and our proposed MapReduce LoRA. Also, Figs. 17 and 18 demonstrate the visual performance of merging progress that advances the visual quality, motion quality, and prompt alignment.
Figure 16: Visual comparison of HunyuanVideo [16], Rewarded Soup [33], and MapReduce LoRA (Ours). In the upper case, MapReduce LoRA more faithfully renders the intended motion and writing action described in the prompt. In the lower case, MapReduce LoRA better adheres to the specified drone-movement control, producing a trajectory that aligns with the prompt.